OpenStudy (nirmalnema):

the product of two numbers is 56 when their sum is added to the sum of their squares we get 128. find the numbers.

OpenStudy (nirmalnema):

plz....reply with whole equation..

OpenStudy (ajprincess):

Let the numbers be a and b. a*b=56 (a+b)+(a^2+b^2)=128 (a+b)^2-2ab=a^2+b^2 (a+b)+(a+b)^2-2ab=128 (a+b)+(a+b)^2-2*56=128 (a+b)+(a+b)^2-112=128 (a+b)+(a+b)^2=128+112 (a+b)+(a+b)^2=240 Assume (a+b)=x x+x^2=240 use quadratic equation and find x

OpenStudy (nirmalnema):

thanks.....

OpenStudy (ajprincess):

have u found x

OpenStudy (nirmalnema):

yup i got it .....:)